OpenStudy (anonymous):

find all real zeros of this polynomial x^3-2x^2-9x+18

OpenStudy (zehanz):

First, try to factor completely. Is is best to separate in two parts: x³-2x²-9x+18 = (x³-2x²)-(9x-18) = x²(x-2) -9(x-2). Can you see now how to factor out one more common factor?

OpenStudy (johnweldon1993):

so you can group these together (x^3 - 3x^2)(-9x+18) what can come out of each? x^2(x - 3) -9(x -2) can you keep going?

OpenStudy (anonymous):

an 8 can come out of each

OpenStudy (anonymous):

x*

OpenStudy (johnweldon1993):

and sorry for my typo inthe first part should be (x^3 - 2x^2)(-9x+18) becomes x^2(x - 2) -9(x -2)

OpenStudy (zehanz):

Once you have x²(x-2) -9(x-2), you can see that x-2 is common, so it can be written as: (x-2)(x²-9). But: x²-9 can itself be further factored. It is of the form a²-b²=(a+b)(a-b), so it is (x+3)(x-3) Putting everything together, we have: (x-2)(x+3)(x-3). To get the zeroes, solve: (x-2)(x+3)(x-3)=0. This is easy now...

OpenStudy (anonymous):

okay that makes sense

OpenStudy (zehanz):

So there are your three real zeroes!

OpenStudy (anonymous):

thanks so much

OpenStudy (zehanz):

YW!